Parth (parthkohli):

A common technique for finding modulo \(10\) of any huge number in the form \(a^b\)?

Parth (parthkohli):

It seems that you have to go through a lot of manipulations... a lot for doing that. Although I was thinking of a technique they use to figure all that out.

Parth (parthkohli):

OK, I'm gonna change the topic to something else:\[39^{39} \equiv n\pmod{100}\]

Parth (parthkohli):

\(n\) has to be a two-digit number. (I actually have to find the last two digits)

Parth (parthkohli):

Something with a pattern?

OpenStudy (shubhamsrg):

39 = -61 mod100 =>39^38 = 61^38 mod100 =>39^39 = 39*61^38 mod100 SInce 61^n always ends in 1, that multiplied by 39, will give last digit as 9, hmm, leme see for the second last.

terenzreignz (terenzreignz):

Better yet, 39^5 = -1 mod 100 Use that: 39^39 = (39^5)^7 * 39^4 = (-1)^7 * 39^4 mod 100

OpenStudy (shubhamsrg):

Seems legit. But on paper, who's gonna go till 39^5 ! :O

Parth (parthkohli):

Very nice, can you tell me *how* you came up with that solution? What do you initially think?

terenzreignz (terenzreignz):

square 39, and then only take the last two digits...

terenzreignz (terenzreignz):

then multiply 39 to it again.

terenzreignz (terenzreignz):

and only take the last two digits, again...

Parth (parthkohli):

Ahem, that was beautiful. But is there a common way to figure out these solutions?

OpenStudy (shubhamsrg):

Ohh, I see.

terenzreignz (terenzreignz):

The way I think of things, and it's a common trend that the way I do it is highly inefficient (LOL) But when faced with a problem b^p = n (mod k) I try to see if there's an exponent to which b can be raised so that it is 1 (or -1) mod k first.

terenzreignz (terenzreignz):

That "last two digits" trick only works because you were conveniently working under mod 100... XD

Parth (parthkohli):

Nice... can you give me another example?

terenzreignz (terenzreignz):

How about 6^5000 = n (mod 215)

terenzreignz (terenzreignz):

(sneakily thinking of simple problems) :>

Parth (parthkohli):

Arhmegurd.\[\left(6^{5}\right)^{1000} \equiv (6^2)^{1000} \equiv n \pmod{215}\]

Parth (parthkohli):

\[\left(6^5\right)^{400} \equiv (6^2)^{400} \equiv n\pmod{215}\]

Parth (parthkohli):

\[\left(6^5\right)^{120} \equiv 6^{240} \equiv n \pmod{215}\]

terenzreignz (terenzreignz):

I'm not following o.O You may be using a method I'm not familiar with, though... if you are, do continue :)

Parth (parthkohli):

\[\left(6^{5}\right)^{48} \equiv 6^{96} \equiv n \pmod{215}\]

Parth (parthkohli):

Hmm... I am just using \(6^2 \equiv 6^5 \pmod{215}\)

Parth (parthkohli):

What should I do here?

terenzreignz (terenzreignz):

Hang on...

terenzreignz (terenzreignz):

I was rather hoping you'd go for \[6^3 = 216 = -1(mod \ 215)\]

Parth (parthkohli):

Oh...

terenzreignz (terenzreignz):

Wait, typo -.-

Parth (parthkohli):

Yes, isn't that 1 (mod 215)

terenzreignz (terenzreignz):

\[\large 6^3 = 216 = 1(\mod \ 215 \ \ \ \ )\] Much better

terenzreignz (terenzreignz):

And work from there...

terenzreignz (terenzreignz):

\[\large 6^{5000} = 6^{3(1666)+2}\]

Parth (parthkohli):

\[6^{5000}= 6^{4998} \times 36 \equiv \]

terenzreignz (terenzreignz):

\[1 \times 36 (\mod 215 \ \ \ \ \ )\]

Parth (parthkohli):

\[\equiv 36 \pmod{215}\]

Parth (parthkohli):

Argh, you type fast. =3

terenzreignz (terenzreignz):

Awesomeness

Parth (parthkohli):

Another n00bish question coming up. Thanks man.

terenzreignz (terenzreignz):

If you have time, look up Fermat's Little Theorem, might come in handy in a pinch.

Parth (parthkohli):

Yesh, that one is cool:\[p^{n - 1} \equiv 1 \pmod{n} \]I'd try to see if that applies in the future.